# Is Your Imitation Learning Policy Better than Mine? Policy Comparison with Near-Optimal Stopping

### v2

@@ Abstract
Imitation learning has enabled robots to perform complex, long-horizon tasks in challenging dexterous manipulation settings. As new methods are developed, they must be rigorously evaluated and compared against corresponding baselines through repeated evaluation trials. However, policy comparison is fundamentally constrained by a small feasible sample size (e.g., 10 or 50) due to significant human effort and limited inference throughput of policies. This paper proposes a novel statistical framework for rigorously comparing two policies in the small sample size regime. Prior work in statistical policy comparison relies on batch testing, which requires a fixed, pre-determined number of trials and lacks flexibility in adapting the sample size to the observed evaluation data. Furthermore, extending the test with additional trials risks inducing inadvertent p-hacking, undermining statistical assurances. In contrast, our proposed statistical test is sequential, allowing researchers to decide whether or not to run more trials based on intermediate results. This adaptively tailors the number of trials to the difficulty of the underlying comparison, saving significant time and effort without sacrificing probabilistic correctness. Extensive numerical simulation and real-world robot manipulation experiments show that our test achieves near-optimal stopping, letting researchers stop evaluation and make a decision in a near-minimal number of trials. Specifically, it reduces the number of evaluation trials by up to 37% as compared to state-of-the-art baselines, while preserving the probabilistic correctness and statistical power of the comparison. Moreover, our method is strongest in the most challenging comparison instances (requiring the most evaluation trials); in a multi-task comparison scenario, we save the evaluator more than 200 simulation rollouts.
@@ I Introduction
To address these challenges, we propose a novel sequential testing framework named STEP (Sequential Testing for Efficient Policy Comparison) for rigorously comparing performance of imitation learning policies [^1] Although this paper focuses on imitation learning, STEP is naturally applicable to evaluating any types of policies based on binary metrics, including reinforcement learning (RL) policies with sparse 0/1 reward.. Unlike batch testing, our approach allows the number of trials to be varied within a given experimental budget. This critical feature offers two practical advantages. First, the evaluator can stop conducting trials early without sacrificing the probabilistic correctness of the comparison if enough statistical evidence is accumulated quickly in favor of $\pi_{1}$ (or $\pi_{0}$). Second, it reduces the epistemic risk of overconfident (and potentially incorrect) conclusions when $\pi_{0}$ and $\pi_{1}$ are closely matched, since the test will abstain from making a decision if statistical evidence remains low after all the planned trials have been executed. Alternatively, the evaluator may safely append additional trials to the original samples and re-conduct statistical analysis on all the results obtained thus far without inadvertent p-hacking. We demonstrate these advantages through simulation and real-world robot manipulation experiments. Furthermore, extensive numerical experiments show that STEP significantly outperforms state-of-the-art (SOTA) sequential methods, reducing the required number of trials by up to 37% without sacrificing probabilistic correctness. The specific contributions of this paper are as follows:
@@ V-A Baseline Procedures
The baselines constitute the SOTA sequential analysis methods described in Lai [30] and Lai and Zhang [31] (termed “Lai”) and the Safe, Anytime-Valid Inference (SAVI) method of Turner and Grünwald [50], which is specifically tailored to contingency table (i.e., policy comparison) problems (termed “SAVI”). Similar to STEP, Lai is a valid sequential method under a pre-determined $N_{\max}$ and is asymptotically optimal as $N_{\max}$ tends to infinity. The specific difference between these methods is that Lai makes the asymptotic assumption of normality of the empirical means. This allows for any distribution subject to the central limit theorem (CLT) to be analyzed tractably in the large-data regime, as the resulting optimal stopping PDE amounts to the solution of a reverse heat equation. STEP, by contrast, solves the exact PDE at the fundamental finite-sample resolution of the testing problem. As will be shown, the Lai procedure does best in difficult (i.e., requiring a lot of data) and symmetric (i.e., $p_{0},p_{1}$ each near $0.5$) test settings, as these best match the asymptotic approximation where the CLT approximation is well-matched to the exact distribution of the empirical means. SAVI, by contrast, does not require a fixed $N_{\max}$ a priori and maintains validity for an arbitrarily large $N_{\max}$. In essence, SAVI procedures construct a test statistic subject to stochastic dynamics (the sequential incorporation of data) which possesses a high-probability ‘stability certificate’ under any null data generating process. The ‘stability’ in this case grants time-uniformity (in perpetuity, i.e., not requiring any finite $N_{\text{max}}$); thus, under any null hypothesis, the probability of ever violating the certificate is bounded (by design) to be less than $\alpha$. In exchange for increased generality and robustness, however, SAVI methods inevitably sacrifice sample complexity in cases where a finite upper-bound to $N_{\max}$ is determined by practical time and resource constraints that researchers are subject to. We emphasize that each existing method is widely applicable to practical testing problems beyond robotics, as is STEP. Importantly, STEP is tailored specifically to the small-$N_{\text{max}}$ evaluation regime.
@@ V-C1 Hardware Evaluation in the Large/Medium-Gap Regime
In FoldRedTowel, we compare two policy checkpoints from a single training run. The baseline policy $\pi_{0}$ was trained for 10000 gradient steps with an AdamW [34] optimizer and achieved the action mean-squared error (MSE) of 1.61E-3 on the validation set. The other policy $\pi_{1}$ continued the training for 70000 additional steps, yielding the validation action MSE of 1.35E-3. To evaluate each policy, five in-distribution (ID) initial conditions were chosen and repeated 10 times each, constituting 50 total trials. ID initial conditions visually resemble the ones that appear in the training dataset. As shown in Table I (rows 1-3), the empirical gap in success rates was 36 percentage points (56% to 92% success), suggesting that $\pi_{0}$ was under-trained. Each sequential method detected a significant difference at level $\alpha^{*}=0.05$ in 19 – 23 trials (apiece). That is, had the algorithms been active during collection, the last 27 – 31 rollouts per policy were unnecessary for confirming the improvement of the later checkpoint over the former. Additionally, note that the Lai method and our STEP sequential procedures were each tuned for an $N_{\text{max}}$ of 50 (row 1), 200 (row 2), and 500 (row 3) rollouts. In the latter two cases, additional rollouts could have been run up to 200 or 500 per policy if the gap was smaller without compromising the validity of the decision.
@@
CleanUpSpill $\mathcal{D}^{i}_{s_{0},o_{0}}$ 0.05 50 50 0.280 0.800 7 9 10 7
CleanUpSpill $\mathcal{D}^{i}_{s_{0},o_{0}}$ 0.05 200 50 0.280 0.800 7 12 10 7
CleanUpSpill $\mathcal{D}^{i}_{s_{0},o_{0}}$ 0.05 500 50 0.280 0.800 7 14 14 7
@@
SpoonOnTowel $\pi_{i}$ 0.01 500 500 0.084 0.386 33 35 33 26
EggplantInBasket $\pi_{i}$ 0.01 500 500 0.400 0.564 192 123 119 128
StackCube $\pi_{i}$ 0.01 500 500 0.000 0.030 329 403 199 135
Multitask $\pi_{i}$ 0.03 1500 1500 N/A N/A 554 561 351 289
@@
TABLE I: Empirical time-to-correct-decision for all hardware (top) and simulation (bottom) policy comparisons. The comparison type is described first; $\pi_{i}$ is comparing two policies, $\mathcal{D}^{i}_{s_{0},o_{0}}$ compares one policy under possible distribution shift. The utilized Type-I Error $\alpha^{*}$ and $N_{\text{max}}$ describe the constraints applied a priori by the evaluator (we underline to emphasize the change in $N_{\text{max}}$ for rows 1-3 and 4-6; observe that the sensitivity of the stopping times is very small). $N~{}\leq~{}N_{\text{max}}$ represents the amount of data available for the statistical analysis. We report the terminal empirical success rates (after $N$ trials) of each policy in each setting under $\hat{p}_{i}$ (this information is not available to any feasible algorithm). We do not have truth labels on this data; however, in all cases, every method arrived at the same decision, including the Oracle SPRT which has a priori access to privileged information $(\hat{p}_{0},\hat{p}_{1})_{N}$; this decision was Reject Null for all rows except the CarrotOnPlate tasks, which each returned Fail To Decide. We report the stopping times of all methods on the right of the table for every context; in all cases: lower is better. We put in bold any feasible method result that is near-optimal within ten trials (absolute) or 25% (relative) of the SPRT Oracle, which is not implementable by an evaluator. In the Multitask setting, we test $p_{1}>p_{0}$ uniformly across the preceding three tasks. This stopping time is the sum by column of the stopping times for the three tasks. Our method saves the evaluator over 200 trials in uniform certification over these three tasks as compared to either feasible baseline.
@@
In CleanUpSpill, we compare the same policy on two different sets of initial conditions. The task is similar to the one originally presented by Xu et al. [60], which compares a set of ID initial conditions against the out-of-distribution (OOD) initial conditions. The ID set includes 10 initial conditions with a white towel and a short blue mug whereas the OOD set uses 10 with a checkered towel with a tall cyan mug (each initial condition is repeated five times). As shown in Table I (rows 4–6), the empirical gap of 52 percentage points was detected in 7–14 trials by all methods, though they were tuned (where applicable) to an $N_{\text{max}}$ up to thirty to seventy times larger; this demonstrates the significant reduction in sensitivity (from an evaluator’s standpoint) arising from setting $N_{\text{max}}$ versus choosing a batch size $N$. Furthermore, STEP’s efficiency only minimally degrades when $N_{\text{max}}$ is increased from 200 (row5) to 500 (row 6). In this setting, any of the three sequential methods would have prevented the need for nearly 70 of the 100 total rollouts (35 of the 50 batch trials per policy).
@@ V-C3 Multi-Task Evaluation in SimplerEnv Simulation
Many established and sophisticated methods exist to efficiently run multi-hypothesis testing (in this case, we are essentially evaluating three separate hypotheses, one for each task [^7] Note that multi-hypothesis testing can naturally handle the case of multi-policy comparison as well, where we would reduce the test to a set of pairwise comparisons that are examined simultaneously.). As a simple illustration, we use the standard Bonferroni (union bound) correction [15] to evaluate the test: specifically, running each of the individual three tests at level $\alpha^{*}_{[\tau]}=0.01$, we observe the stopping times shown in Table I, rows 9-11 (right-hand side). Via the Bonferroni correction, the combined decision (RejectNull, because every subtest decided RejectNull) expressed in LABEL:Eqn:MultiHypothesis is then confirmed at $\alpha=0.03=\sum_{\tau}\alpha^{*}_{[\tau]}$. As illustrated in Table I, each sequential method saves a substantial number of simulation rollouts on the easiest subtest (SpoonOnTowel). As expected, SAVI begins to struggle when the tests become more challenging, such as in EggplantInBasket. Finally, we observe the weakness of the Lai procedure: in heavily skewed cases, it suffers substantially even compared with SAVI methods, as shown in StackCube. To summarize: naive multitask evaluation requires the aggregation of multiple batches of rollouts, here totaling 1500 per policy (500 per task per policy). Note in rows 9-11 how different the number of requisite trials can be, and therefore how hard it is to reliably run the evaluation using a fixed batch size. On the easiest task, even when tuned to $N_{\text{max}}=500$, the comparison was answered in fewer than 40 rollouts by all sequential methods, a savings of over 90%. On the progressively harder cases the number of required samples increased 5-10 times over the easiest, but our method (STEP) improved substantially over each of the other sequential procedures. In total, STEP would have saved the evaluator an additional 200 rollouts for each of Octo-small and Octo-base for the multitask comparison problem as compared to the current SOTA approaches.
@@ Acknowledgments
This work was primarily completed during an internship at Toyota Research Institute (TRI). The authors would like to thank Chen Xu for insightful feedback and discussions, as well as the robot teacher team at TRI headquarters for their data collection efforts, in particular Derick Seale and Donovan Jackson. Finally, we thank the anonymous reviewers and the area chair for incisive suggestions for improving the paper.

### v4

@@ Abstract
Imitation learning has enabled robots to perform complex, long-horizon tasks in challenging dexterous manipulation settings. As new methods are developed, they must be rigorously evaluated and compared against corresponding baselines through repeated evaluation trials. However, policy comparison is fundamentally constrained by a small feasible sample size (e.g., 10 or 50) due to significant human effort and limited inference throughput of policies. This paper proposes a novel statistical framework for rigorously comparing two policies in the small sample size regime. Prior work in statistical policy comparison relies on batch testing, which requires a fixed, pre-determined number of trials and lacks flexibility in adapting the sample size to the observed evaluation data. Furthermore, extending the test with additional trials risks inducing inadvertent p-hacking, undermining statistical assurances. In contrast, our proposed statistical test is sequential, allowing researchers to decide whether or not to run more trials based on intermediate results. This adaptively tailors the number of trials to the difficulty of the underlying comparison, saving significant time and effort without sacrificing probabilistic correctness. Extensive numerical simulation and real-world robot manipulation experiments show that our test achieves near-optimal stopping, letting researchers stop evaluation and make a decision in a near-minimal number of trials. Specifically, it reduces the number of evaluation trials by up to 32% as compared to state-of-the-art baselines, while preserving the probabilistic correctness and statistical power of the comparison. Moreover, our method is strongest in the most challenging comparison instances (requiring the most evaluation trials); in a multi-task comparison scenario, we save the evaluator more than 160 simulation rollouts.
@@ I Introduction
To address these challenges, we propose a novel sequential testing framework named STEP (Sequential Testing for Efficient Policy Comparison) for rigorously comparing performance of imitation learning policies [^1] Although this paper focuses on imitation learning, STEP is naturally applicable to evaluating any types of policies based on binary metrics, including reinforcement learning (RL) policies with sparse 0/1 reward.. Unlike batch testing, our approach allows the number of trials to be varied within a given experimental budget. This critical feature offers two practical advantages. First, the evaluator can stop conducting trials early without sacrificing the probabilistic correctness of the comparison if enough statistical evidence is accumulated quickly in favor of $\pi_{1}$ (or $\pi_{0}$). Second, it reduces the epistemic risk of overconfident (and potentially incorrect) conclusions when $\pi_{0}$ and $\pi_{1}$ are closely matched, since the test will abstain from making a decision if statistical evidence remains low after all the planned trials have been executed. Alternatively, the evaluator may safely append additional trials to the original samples and re-conduct statistical analysis on all the results obtained thus far without inadvertent p-hacking. We demonstrate these advantages through simulation and real-world robot manipulation experiments. Furthermore, extensive numerical experiments show that STEP significantly outperforms state-of-the-art (SOTA) sequential methods, reducing the required number of trials by up to 32% without sacrificing probabilistic correctness. The specific contributions of this paper are as follows:
@@ V-A Baseline Procedures
The baselines constitute the SOTA sequential analysis methods described in Lai [30] and Lai and Zhang [31] (termed “Lai”) and the Safe, Anytime-Valid Inference (SAVI) method of Turner and Grünwald [50], which is specifically tailored to contingency table (i.e., policy comparison) problems (termed “SAVI”).
Similar to STEP, Lai is a valid sequential method under a pre-determined $N_{\max}$ and is asymptotically optimal as $N_{\max}$ tends to infinity. The specific difference between these methods is that Lai makes the asymptotic assumption of normality of the empirical means. This allows for any distribution subject to the central limit theorem (CLT) to be analyzed tractably in the large-data regime, as the resulting optimal stopping PDE amounts to the solution of a reverse heat equation. STEP, by contrast, solves the exact PDE at the fundamental finite-sample resolution of the testing problem. As will be shown, the Lai procedure does best in difficult (i.e., requiring a lot of data) and symmetric (i.e., $p_{0},p_{1}$ each near $0.5$) test settings, as these best match the asymptotic approximation where the CLT approximation is well-matched to the exact distribution of the empirical means. SAVI, by contrast, does not require a fixed $N_{\max}$ a priori and maintains validity for an arbitrarily large $N_{\max}$. In essence, SAVI procedures construct a test statistic subject to stochastic dynamics (the sequential incorporation of data) which possesses a high-probability ‘stability certificate’ under any null data generating process. The ‘stability’ in this case grants time-uniformity (in perpetuity, i.e., not requiring any finite $N_{\text{max}}$); thus, under any null hypothesis, the probability of ever violating the certificate is bounded (by design) to be less than $\alpha$. In exchange for increased generality and robustness, however, SAVI methods inevitably sacrifice sample complexity in cases where a finite upper-bound to $N_{\max}$ is determined by practical time and resource constraints that researchers are subject to. We emphasize that each existing method is widely applicable to practical testing problems beyond robotics, as is STEP. Importantly, STEP is tailored specifically to the small-$N_{\text{max}}$ evaluation regime.
@@ V-C1 Hardware Evaluation in the Large/Medium-Gap Regime
In FoldRedTowel, we compare two policy checkpoints from a single training run. The baseline policy $\pi_{0}$ was trained for 10000 gradient steps with an AdamW [34] optimizer and achieved the action mean-squared error (MSE) of 1.61E-3 on the validation set. The other policy $\pi_{1}$ continued the training for 70000 additional steps, yielding the validation action MSE of 1.35E-3. To evaluate each policy, five in-distribution (ID) initial conditions were chosen and repeated 10 times each, constituting 50 total trials. ID initial conditions visually resemble the ones that appear in the training dataset. As shown in Table I (rows 1-3), the empirical gap in success rates was 36 percentage points (56% to 92% success), suggesting that $\pi_{0}$ was under-trained. Each sequential method detected a significant difference at level $\alpha^{*}=0.05$ in 17 – 23 trials (apiece). That is, had the algorithms been active during collection, the last 27 – 33 rollouts per policy were unnecessary for confirming the improvement of the later checkpoint over the former. Additionally, note that the Lai method and our STEP sequential procedures were each tuned for an $N_{\text{max}}$ of 50 (row 1), 200 (row 2), and 500 (row 3) rollouts. In the latter two cases, additional rollouts could have been run up to 200 or 500 per policy if the gap was smaller without compromising the validity of the decision.
@@
CleanUpSpill $\mathcal{D}^{i}_{s_{0},o_{0}}$ 0.05 50 50 0.280 0.800 7 8 8 7
CleanUpSpill $\mathcal{D}^{i}_{s_{0},o_{0}}$ 0.05 200 50 0.280 0.800 7 13 9 7
CleanUpSpill $\mathcal{D}^{i}_{s_{0},o_{0}}$ 0.05 500 50 0.280 0.800 7 14 13 7
@@
SpoonOnTowel $\pi_{i}$ 0.01 500 500 0.084 0.386 33 36 36 26
EggplantInBasket $\pi_{i}$ 0.01 500 500 0.400 0.564 192 125 131 128
StackCube $\pi_{i}$ 0.01 500 500 0.000 0.030 329 417 225 135
Multitask $\pi_{i}$ 0.03 1500 1500 N/A N/A 554 578 392 289
@@
TABLE I: Empirical time-to-correct-decision for all hardware (top) and simulation (bottom) policy comparisons. The comparison type is described first; $\pi_{i}$ is comparing two policies, $\mathcal{D}^{i}_{s_{0},o_{0}}$ compares one policy under possible distribution shift. The utilized Type-I Error $\alpha^{*}$ and $N_{\text{max}}$ describe the constraints applied a priori by the evaluator (we underline to emphasize the change in $N_{\text{max}}$ for rows 1-3 and 4-6; observe that the sensitivity of the stopping times is very small). $N~{}\leq~{}N_{\text{max}}$ represents the amount of data available for the statistical analysis. We report the terminal empirical success rates (after $N$ trials) of each policy in each setting under $\hat{p}_{i}$ (this information is not available to any feasible algorithm). We do not have truth labels on this data; however, in all cases, every method arrived at the same decision, including the Oracle SPRT which has a priori access to privileged information $(\hat{p}_{0},\hat{p}_{1})_{N}$; this decision was Reject Null for all rows except the CarrotOnPlate tasks, which each returned Fail To Decide. We report the stopping times of all methods on the right of the table for every context; in all cases: lower is better. We put in bold any feasible method result that is near-optimal within ten trials (absolute) or 25% (relative) of the SPRT Oracle, which is not implementable by an evaluator. In the Multitask setting, we test $p_{1}>p_{0}$ uniformly across the preceding three tasks. This stopping time is the sum by column of the stopping times for the three tasks. Our method saves the evaluator over 160 trials in uniform certification over these three tasks as compared to either feasible baseline.
@@
In CleanUpSpill, we compare the same policy on two different sets of initial conditions. The task is similar to the one originally presented by Xu et al. [60], which compares a set of ID initial conditions against the out-of-distribution (OOD) initial conditions. The ID set includes 10 initial conditions with a white towel and a short blue mug whereas the OOD set uses 10 with a checkered towel with a tall cyan mug (each initial condition is repeated five times). As shown in Table I (rows 4–6), the empirical gap of 52 percentage points was detected in 7–14 trials by all methods, though they were tuned (where applicable) to an $N_{\text{max}}$ up to thirty to seventy times larger; this demonstrates the significant reduction in sensitivity (from an evaluator’s standpoint) arising from setting $N_{\text{max}}$ versus choosing a batch size $N$. Furthermore, STEP’s efficiency only minimally degrades when $N_{\text{max}}$ is increased from 200 (row 5) to 500 (row 6). In this setting, any of the three sequential methods would have prevented the need for at least 70 of the 100 total rollouts (35 of the 50 batch trials per policy).
@@ V-C3 Multi-Task Evaluation in SimplerEnv Simulation
Many established and sophisticated methods exist to efficiently run multi-hypothesis testing (in this case, we are essentially evaluating three separate hypotheses, one for each task [^7] Note that multi-hypothesis testing can naturally handle the case of multi-policy comparison as well, where we would reduce the test to a set of pairwise comparisons that are examined simultaneously.). As a simple illustration, we use the standard Bonferroni (union bound) correction [15] to evaluate the test: specifically, running each of the individual three tests at level $\alpha^{*}_{[\tau]}=0.01$, we observe the stopping times shown in Table I, rows 9-11 (right-hand side). Via the Bonferroni correction, the combined decision (RejectNull, because every subtest decided RejectNull) expressed in LABEL:Eqn:MultiHypothesis is then confirmed at $\alpha=0.03=\sum_{\tau}\alpha^{*}_{[\tau]}$. As illustrated in Table I, each sequential method saves a substantial number of simulation rollouts on the easiest subtest (SpoonOnTowel). As expected, SAVI begins to struggle when the tests become more challenging, such as in EggplantInBasket. Finally, we observe the weakness of the Lai procedure: in heavily skewed cases, it suffers substantially even compared with SAVI methods, as shown in StackCube. To summarize: naive multitask evaluation requires the aggregation of multiple batches of rollouts, here totaling 1500 per policy (500 per task per policy). Note in rows 9-11 how different the number of requisite trials can be, and therefore how hard it is to reliably run the evaluation using a fixed batch size. On the easiest task, even when tuned to $N_{\text{max}}=500$, the comparison was answered in fewer than 40 rollouts by all sequential methods, a savings of over 90%. On the progressively harder cases the number of required samples increased 5-10 times over the easiest, but our method (STEP) improved substantially over each of the other sequential procedures. In total, STEP would have saved the evaluator an additional 160 rollouts for each of Octo-small and Octo-base for the multitask comparison problem as compared to the current SOTA approaches.
@@ Acknowledgments
This work was primarily completed during an internship at Toyota Research Institute (TRI). Anirudha Majumdar was partially supported by the NSF CAREER Award [#2044149] and the Sloan Fellowship. The authors would like to thank Chen Xu for insightful feedback and discussions, as well as the robot teacher team at TRI headquarters for their data collection efforts, in particular Derick Seale and Donovan Jackson. Finally, we thank the anonymous reviewers and the area chair for incisive suggestions for improving the paper.